Question: Problem 8-4 (Algorithmic) Modified Accelerated Cost Recovery System (MACRS) (LO 8.2) On March 8, ...

Problem 8-4 (Algorithmic) Modified Accelerated Cost Recovery System (MACRS) (LO 8.2) On March 8, 2018, Holly purchased a residential apartment building. The cost basis assigned to the building is $195,400. Holly also owns another residential apartment building that she purchased on December 15, 2018, with a cost basis of $790,000. Click here to access the depreciation tables. If required, round intermediate calculations and final answers to nearest dollar.

a. Calculate Holly's total depreciation deduction for the apartments for 2018 using MACRS.

b. Calculate Holly's total depreciation deduction for the apartments for 2019 using MACRS.

TABLE 8.5
Straight-Line Depreciation for Real Property Assuming Mid-Month Convention*
27.5-Year Residential Real Property
The applicable annual percentage is (use the column for the
month in the first year the property is placed in service):
Recovery
Year(s) 1 2 3 4 5 6 7 8 9 10 11 12
1 3.485 3.182 2.879 2.576 2.273 1.970 1.667 1.364 1.061 0.758 0.455 0.152
2–18 3.636 3.636 3.636 3.636 3.636 3.636 3.636 3.636 3.636 3.636 3.636 3.636
19–27 3.637 3.637 3.637 3.637 3.637 3.637 3.637 3.637 3.637 3.637 3.637 3.637
28 1.970 2.273 2.576 2.879 3.182 3.485 3.636 3.636 3.636 3.636 3.636 3.636
29 0.000 0.000 0.000 0.000 0.000 0.000 0.152 0.455 0.758 1.061 1.364 1.667
39-Year Nonresidential Real Property
The applicable annual percentage is (use the column for the
month in the first year the property is placed in service):
Recovery
Year(s)
1 2 3 4 5 6 7 8 9 10 11 12
1 2.461 2.247 2.033 1.819 1.605 1.391 1.177 0.963 0.749 0.535 0.321 0.107
2–39 2.564 2.564 2.564 2.564 2.564 2.564 2.564 2.564 2.564 2.564 2.564 2.564
40 0.107 0.321 0.535 0.749 0.963 1.177 1.391 1.605 1.819 2.033 2.247 2.461
*The official tables contain a separate row for each year. For ease of presentation, certain years are grouped together in these two tables. In some instances, this will produce a difference of .001 percent when compared with the official tables.
Answer #1

Part A

Holly's total depreciation deduction for the apartments for 2018 = (195400*2.879%)+(790000*0.152%) = $6826

Part B

Holly's total depreciation deduction for the apartments for 2019 = (195400*3.636%)+(790000*3.636%) = $35829